# fixedprepend, caption=FIXED, color=green

### v2

@@ 6 Conclusion
In this work, we have developed a chemical reasoner LLM, ChemDFM-R, by tackling both the limitations in understanding atomized chemical knowledge and the domain-specific reasoning logic. By incorporating atomized knowledge about molecular functional groups and their changes during reactions into the pretraining corpus, and applying a mix-sourced distillation approach before reinforcement learning, we have enhanced the model’s ability to reason efficiently and effectively in chemistry. Our extensive experiments demonstrate that ChemDFM-R significantly improves chemical problem-solving and reasoning capabilities, making it a valuable tool for facilitating reliable human-AI collaboration in chemistry and advancing AI-driven research and applications.
Appendix A Functional Groups covered by ChemFG
@@ B.1.1 Molecule-Centric Tasks
  • Property Classification: These tasks require models to predict the value of molecular properties from a list of candidates (usually yes and no). The data is constructed from 5 of the most popular property classification datasets in MoleculeNet [Wu et al., 2018], namely BACE, BBBP, ClinTox, HIV, and Tox21.
@@ B.1.2 Reaction-Centric Tasks
  • Reaction Completion: Given an incomplete reaction, models need to complete the missing reactants, reagents, or products. Raw data comes from USPTO-Full [Dai et al., 2019], USPTO-MIT [Jin et al., 2017], and USPTO-50K [Schneider et al., 2016].
  • Step Prediction: Given a reaction, models are required to predict the experimental procedure to conduct it in the laboratory. Raw data comes from USPTO [Dai et al., 2019].
@@ B.1.3 Knowledge-Centric Tasks
  • Literature QA: In this task, models are required to answer questions based on the given paragraph. The data is extracted from the long paragraph following the method in SciQAG [Wan et al., 2024]. The raw data comes from the articles in the domain-pertaining. The articles are split into sections and then truncated into paragraphs within 2k to 3k tokens. We ask GPT-4o-mini to extract 15 keywords from each paragraph, then generate 10 question-answer pairs according to them. We adopt another LLM, Qwen2.5-14B-Instruct, to evaluate the quality of the QA pair in 4 dimensions: completeness, accuracy, reasonableness, and agnosticism. The LLM will score the QA pair from 1 to 5 using the designed prompts. QA pairs with any scores below 5 are discarded. If there are more than 1 QA pair left, the questions are asked in conversation turns.
@@ B.2 Instruction Generation
Appendix C Comparison of Rationales Generated by Teacher Models
@@
As shown in Figure 28, and 9, the rationales generated using our method exhibit distinct advantages in terms of accuracy, completeness, and logicality.
@@ Appendix D Details of Reinforcement Learning
We use the DAPO [Yu et al., 2025] algorithm for reinforcement learning. For each training sample, we sample 8 different rollouts during training. The learning rate is set to 5e-7, and the batch sizes for rollouts and training are set to 512 and 128, respectively. As for the reward system, we follow the method recommended by DeepSeek-R1 [Guo et al., 2025], where the reward system consists of format rewards and accuracy rewards. The format rewards evaluate whether a response strictly follows the reasoning format, and the accuracy rewards evaluate whether a response is correct. Specifically, considering the redundancy of the casual SMILES notations (one molecule could correspond to multiple SMILES), we first canonicalize all the SMILES in the response before calculating the accuracy rewards.
Appendix E Detail Results of Benchmark Evaluation
E.1 ChemEval
We consider the L1 and L2 level tasks in ChemEval as text-centric tasks, the L3 level tasks as molecule-centric tasks, and the L4 level tasks as reaction-centric tasks. Moreover, there are tasks that we can not achieve a feasible grading in ChemEval. We temporarily skip these tasks. The raw results are demonstrated in Table 3.
Tasks Metric MolInst ChemLLM ChemDFM ChemDFM GPT-4o Qwen3-14B DeekSeek Qwen3-14B o4-mini ChemDFM-R -20B-DPO -13B-v1.0 -8B-v1.5 (no think) -R1 (think) Text-Centric Tasks MCTask Acc 0.40 0.35 0.50 0.00 0.75 0.75 0.90 0.80 0.65 0.65 FBTask Acc 0.00 0.00 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TFTask Acc 0.55 0.75 0.70 0.05 0.80 0.85 0.95 0.80 0.85 0.85 SATask BLEU 0.01 0.17 0.12 0.11 0.15 0.09 0.04 0.07 0.10 0.13 CNER F1 0.00 0.44 0.59 0.23 0.67 0.61 0.70 0.67 0.57 0.62 CERC F1 0.05 0.00 0.12 0.00 0.20 0.24 0.28 0.20 0.26 0.26 AddE F1 0.34 0.59 0.47 0.05 0.70 0.65 0.71 0.65 0.68 0.72 SolvE F1 0.58 0.70 0.83 0.25 0.75 0.73 0.79 0.54 0.72 0.73 TempE F1 0.70 0.75 0.20 0.25 0.75 0.95 0.95 0.90 0.95 0.95 TimeE F1 0.95 0.95 0.40 0.95 0.95 0.95 0.95 0.95 0.95 0.95 CharME F1 0.21 0.47 0.26 0.49 0.79 0.74 0.71 0.61 0.76 0.49 CatTE F1 0.80 0.90 0.55 0.30 0.95 0.85 1.00 0.85 0.90 0.75 YieldE F1 0.00 0.55 0.20 0.30 0.80 0.75 0.75 0.95 0.75 0.60 AbsGen BLEU 0.56 0.46 0.54 0.04 0.67 0.70 0.62 0.65 0.67 0.69 OLGen BLEU 0.19 0.12 0.22 0.02 0.14 0.24 0.10 0.18 0.13 0.15 TopC Acc 0.40 0.50 0.40 0.20 0.50 0.40 0.55 0.50 0.50 0.50 ReactTR F1 0.20 0.13 0.15 0.05 0.35 0.30 0.25 0.30 0.20 0.30 Molecule-Centric Tasks MolNG BLEU 0.04 0.57 0.76 0.35 0.60 0.43 0.10 0.42 0.69 0.84 IUPAC2MF EM 0.00 0.00 0.25 0.00 0.40 0.00 0.05 0.00 0.60 0.75 SMILES2MF EM 0.00 0.00 0.45 0.00 0.10 0.05 0.00 0.00 0.55 0.85 IUPAC2SMILES EM 0.00 0.00 0.10 0.00 0.00 0.00 0.05 0.00 0.15 0.55 SMILES2IUPAC EM 0.00 0.00 0.00 0.05 0.00 0.00 0.00 0.00 0.00 0.35 S2S EM 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 MolPC Acc 0.50 0.53 0.61 0.09 0.64 0.54 0.57 0.59 0.66 0.73 Mol2PC BLEU 0.30 0.36 0.39 0.19 0.37 0.41 0.30 0.38 0.34 0.57 Reaction-Centric Tasks SubRec Acc 0.00 0.03 0.08 0.02 0.08 0.00 0.05 0.00 0.18 0.25 LRec Acc 0.00 0.00 0.20 0.00 0.10 0.05 0.10 0.35 0.00 0.45 RRec Acc 0.00 0.00 0.19 0.00 0.10 0.15 0.05 0.25 0.30 0.35 SolvRec Acc 0.00 0.00 0.05 0.00 0.35 0.02 0.25 0.35 0.30 0.40 CatRec Acc 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TempRec RMSE $\downarrow$ 0.77 0.27 0.71 0.22 0.22 0.23 0.18 0.35 0.22 0.30 TimeRec RMSE $\downarrow$ 0.20 0.19 0.18 0.09 0.07 0.08 0.10 0.11 0.10 0.15 PPre Acc 0.00 0.00 0.03 0.05 0.10 0.00 0.18 0.06 0.35 0.60 YPred Acc 0.20 0.20 0.20 0.00 0.55 0.15 0.25 0.25 0.45 0.20 RatePred Overlap 0.03 0.14 0.05 0.09 0.18 0.15 0.31 0.10 0.07 0.09 IMDer Acc 0.00 0.00 0.10 0.00 0.05 0.00 0.05 0.00 0.05 0.00
Table 3: The detailed benchmark results of different models on ChemEval.
As illustrated in Table 3, ChemDFM-R manages to achieve competitive performance in the text-centric tasks compared with the cutting-edge LLMs, while achieve SOTA performance across all the molecule-centric tasks and a large portion of the reaction-centric tasks. A detailed analysis of the task characteristics reveals that ChemDFM-R tends to perform less effectively on tasks involving numerical prediction, which will be a key focus of our future optimization efforts.
E.2 SciKnowEval
We group the tasks in SciKnowEval based on their input and output. Specifically, the task is classified as a text-centric task if there is no SMILES appear in its input or output, as a reaction-centric task if there are reaction SMILES appear in its input or output, and as a molecule-centric task otherwise. Due to budget limit, we currently skip the tasks that require GPT-4o for grading. The raw results are demonstrated in Table 4.
Tasks MolInst ChemLLM ChemDFM ChemDFM GPT-4o Qwen3-14B DeekSeek Qwen3-14B o4-mini ChemDFM-R -20B-DPO -13B-v1.0 -8B-v1.5 (no think) -R1 (think) Text-Centric Tasks Chem LiterQA 0.75 0.79 0.78 0.80 0.86 0.84 0.86 0.84 0.89 0.80 React Mech Infer. 0.96 0.99 0.98 0.99 1.00 0.99 0.97 0.98 0.98 0.98 Comp Iden. and Prop. 0.94 0.97 0.95 0.97 0.99 0.98 0.96 0.99 0.99 0.98 Chem DU 0.95 0.98 0.97 0.97 1.00 0.99 0.99 0.99 0.99 0.98 Chem HV 0.83 0.89 0.86 0.88 0.95 0.94 0.93 0.92 0.93 0.92 Chem RI 0.93 0.96 0.95 0.96 0.97 0.97 0.95 0.96 0.96 0.97 Balancing Eq. 0.01 0.16 0.12 0.22 0.10 0.08 0.11 0.73 0.10 0.27 Chem Cal. 0.35 0.33 0.37 0.37 0.52 0.54 0.79 0.91 0.92 0.54 Mol Tox. Pred 0.54 0.54 0.44 0.45 0.37 0.52 0.62 0.51 0.55 0.44 Chem Safe Test 0.69 0.75 0.60 0.65 0.85 0.81 0.82 0.83 0.80 0.81 Molecule-Centric Tasks Mol Name Conv. 0.60 0.71 0.70 0.83 0.87 0.71 0.42 0.84 0.96 0.90 Mol Prop. Pred 0.32 0.34 0.38 0.48 0.44 0.38 0.41 0.42 0.48 0.45 Mol Cap. 0.14 0.21 0.23 0.22 0.14 0.09 0.13 0.10 0.14 0.51 Mol Weight Cal. 0.26 0.26 0.17 0.23 0.26 0.17 0.04 0.31 0.63 0.37 Mol Prop. Cal. 0.39 0.33 0.31 0.30 0.35 0.37 0.39 0.38 0.45 0.20 Mol Stru. Pred 0.33 0.31 0.35 0.26 0.46 0.31 0.07 0.37 0.57 0.34 Mol Gen. 0.64 0.53 0.84 0.89 0.61 0.39 0.69 0.44 0.75 0.83 Reaction-Centric Tasks Reaction Pred 0.39 0.93 0.92 0.96 0.48 0.91 0.82 0.94 0.99 0.99 Retrosynthesis 0.44 0.83 0.78 0.77 0.41 0.70 0.83 0.83 0.95 0.90
Table 4: The detailed benchmark results of different models on SciKnowEval.
Appendix F More Rationale Analysis

### v3

@@ 6 Conclusion
In this work, we have developed a chemical reasoning LLM, ChemDFM-R, by tackling both the limitations in understanding atomized chemical knowledge and the domain-specific reasoning logic. By incorporating atomized knowledge about molecular functional groups and their changes during reactions into the pretraining corpus, and applying a mixed-source distillation approach before reinforcement learning, we have enhanced the model’s ability to reason efficiently and effectively in chemistry. Our extensive experiments demonstrate that ChemDFM-R significantly improves chemical problem-solving and reasoning capabilities, making it a valuable tool for facilitating reliable human-AI collaboration in chemistry and advancing AI-driven research and applications.
Disclosing of LLM usage
LLMs are leveraged exclusively as an assist tool for paper writing. We only use them to help us polish the language and improve readability. They do not contribute to any of the creative work involved in this paper, such as research ideation, experiment design, data analysis, and result interpretation. Therefore, they should not be regarded as authorship or substantive contribution.
Appendix A Details about ChemFG
A.1 Raw Data Collection
Literature.
Literature, including papers and textbooks, contains not only the widely accepted chemical knowledge and principles, but also the cutting-edge research in the field of chemistry. Therefore, to take full advantage of chemical literature, we collected over 12 million literature from the open Internet dated prior to January 2022. After further cleaning and deduplication, 79B tokens are obtained from it.
Molecules.
Molecules are the fundamental participants in various chemical processes. Therefore, it is crucial for Chemical LLMs to understand molecular structures and properties. We manage to acquire large-scale molecule datasets from PubChem [^5] https://pubchem.ncbi.nlm.nih.gov/, one of the biggest open accessible chemical databases with more than 100M compounds. We include 30 million molecules along with their notations, descriptions (if applicable), and properties. Besides PubChem, we also leverage the PubChemQC [nakata2017pubchemqc] dataset, which contains the quantum chemical calculation results of 86M molecules from PubChem, to supplement the quantum chemical properties of these molecules, such as dipole moment and orbital energy. To diversify the final data entry, we randomize the order of the properties and use three different formats: markdown list, markdown table, and JSON dictionary to formulate the molecule data.
Reactions.
Reactions are the major process in the chemical world. In ChemFG, we use the reactions from USPTO-FULL [dai2019gln], one of the most comprehensive open-sourced chemical reaction databases. To avoid data leakage, we exclude the test set of USPTO-FULL, USPTO-MIT [jin2017predicting], and USPTO-50K [schneider2016s] according to the products of reactions. Moreover, to further enhance the data diversity, we leverage the SMILES (Simplified Molecular Input Line Entry System) augmentation method introduced in R-SMILES [zhong2022rsmiles] and achieve a total of 10 times augmentation of data. Finally, a corpus of 7 million reactions is obtained.
A.2 Functional Groups Coverage
@@
A.3 Quality Control
To validate the correctness of our functional group identification toolkit, we hired three graduate-level chemical experts to conduct manual inspections. Specifically, we sampled 100 annotated molecules and reactions, respectively, and asked the experts to determine whether the annotations were correct. Results show that our functional group identification toolkit achieves 98% accuracy rate on molecules and 89% on reactions. Examples of the errors are demonstrated in Figure 7.
Figure 7: Examples of the error annotations of our functional group identification toolkits.
@@ B.1.1 Molecule-Centric Tasks
  • Property Classification: These tasks require models to predict the value of molecular properties from a list of candidates (usually yes and no). The data is constructed from 5 of the most popular property classification datasets in MoleculeNet [wu2018moleculenet], namely BACE, BBBP, ClinTox, HIV, and Tox21.
@@ B.1.2 Reaction-Centric Tasks
  • Reaction Completion: Given an incomplete reaction, models need to complete the missing reactants, reagents, or products. Raw data comes from USPTO-Full [dai2019gln], USPTO-MIT [jin2017predicting], and USPTO-50K [schneider2016s].
  • Step Prediction: Given a reaction, models are required to predict the experimental procedure to conduct it in the laboratory. Raw data comes from USPTO [dai2019gln].
@@ B.1.3 Knowledge-Centric Tasks
  • Literature QA: In this task, models are required to answer questions based on the given paragraph. The data is extracted from the long paragraph following the method in SciQAG [wan2024sciqagframeworkautogeneratedscience]. The raw data comes from the articles in the domain-pertaining. The articles are split into sections and then truncated into paragraphs within 2k to 3k tokens. We ask GPT-4o-mini to extract 15 keywords from each paragraph, then generate 10 question-answer pairs according to them. We adopt another LLM, Qwen2.5-14B-Instruct, to evaluate the quality of the QA pair in 4 dimensions: completeness, accuracy, reasonableness, and agnosticism. The LLM will score the QA pair from 1 to 5 using the designed prompts. QA pairs with any scores below 5 are discarded. If there are more than 1 QA pair left, the questions are asked in conversation turns.
@@ B.2 Instruction Generation
Appendix C Quality Control of the Distillation Rationale Generation
C.1 Comparison of Rationales Generated by Teacher Models
@@
As shown in Figure 39, and 10, the rationales generated using our method exhibit distinct advantages in terms of accuracy, completeness, and logicality.
@@
C.2 Quality Control
Figure 11: Human validation result of teachers’ rationales. “Minor Flaw” represents acceptable flaws, such as skipping reasoning steps or missing possibilities. “Substantive Issue” means severe logic errors or nonsense reasoning.
To quantitatively validate the quality of the teachers’ rationales generated by our method, we hired three graduate-level chemical experts to perform manual assessments. Results in Figure 11 show that among the sampled 264 rationales, 177 of them (67%) exhibit sufficiently high quality, 60 of them (23%) have minor, acceptable flaws such as reasoning step skipping or missing possibilities, and 27 of them (10%) contain substantive issues such as logic errors or nonsense reasoning. Two examples of distilled rationales with minor flaws are shown in Figure 12, and two examples of distilled rationales with substantive issues are shown in Figure 13. Considering that reinforcement learning will be used after distillation to correct errors and improve performance, and that it is difficult to systematically distinguish these problems in the rationales, we used all the data during the distillation process.
Figure 12: Examples of the distilled rationales with minor flaws.
Figure 13: Examples of the distilled rationales with substantive issues.
@@ Appendix D Details of Reinforcement Learning
To construct the reinforcement learning dataset, a subset of tasks that have verifiable answers is selected from our instruction-tuning dataset. We sampled data from the same sources used to build the instruction-tuning dataset, while minimizing overlap between the specific instances contained in the two datasets. These tasks are mixed together and uniformly sampled during training. We use the DAPO [yu2025dapo] algorithm for reinforcement learning. As for the reward system, we follow the method recommended by DeepSeek-R1 [guo2025deepseek], where the reward system consists of format rewards and accuracy rewards. The format rewards evaluate whether a response strictly follows the reasoning format, and the accuracy rewards evaluate whether a response is correct. Specifically, considering the redundancy of the casual SMILES notations (one molecule could correspond to multiple SMILES), we first canonicalize all the SMILES in the response before calculating the accuracy rewards.
Appendix E Training Details of ChemDFM-R
E.1 Mitigating Data Leakage
Data leakage is a crucial problem for the evaluation of LLMs. During the training of ChemDFM-R, measures have been taken to reduce the risk of data leakage. For the domain pretraining corpus, we avoid using the same molecules and reactions as those presented in SciKnowEval [feng2024sciknoweval] and ChemEval [huang2024chemeval] benchmarks. For instruction tuning and mixed-source distillation, the same molecules and reactions are deliberately excluded from a task if they appear in relative tasks in SciKnowEval and ChemEval. We use canonicalized SMILES to determine whether two molecules are the same. Reactions are considered the same if they share the same product.
E.2 Statistics about Computation and Data
We trained ChemDFM-R on NVIDIA A800 Tensor Core GPUs for a total of 30840 GPU hours. Specifically:
  • 1).
    for domain pretraining, the model was trained on the 101 billion[-] ChemFG corpus A for 24728 GPU hours;
  • 2).
    for instruction tuning, the model was trained on over 7.5 million instructions B, which consist of 2.5 million chemical instructions and 5 million general instructions, for 3785 GPU hours;
  • 3).
    for mixed-sourced distillation, the model was trained on the mixed-sourced distillation dataset, which is of the same scale as the instruction-tuning dataset, for 2059 GPU hours;
  • 4).
    for reinforcement learning, the model was trained on 121,811 samples for 268 GPU hours.
E.3 Hyperparameters and System Prompts
Training Hyperparameters.
The hyperparameters used during the training of ChemDFM-R are listed in Table 6.
Inference Setting.
During the inference of ChemDFM-R, we set the temperature to 0.6, topK to 1, and topP to 1 with no other penalties.
System Prompts.
During the training of ChemDFM-R, three different system prompts are used. Specifically:
  • For all samples in the instruction-tuning dataset and the non-reasoning samples in the mixed-source distillation dataset, we use: "You are a helpful assistant."
[…]
Table 6: The training hyperparameters used during the training of ChemDFM-R.
Domain Instruction Mixed-source Reinforcement
Pretraining Tuning Distillation Learning
Initial Learning Rate 1e-5 1e-5 1e-5 5e-7
Minimal Learning Rate 1e-6 0 0 0
Optimizer Adam(0.9, 0.95)
Scheduler Cosine
Max Sequence Length 8192 8192 8192 -
Max Generation Length - - - 1024
Train Batch Size 624 512 512 128
Rollout Batch Size - - - 512
Epochs 1 2 1 1
DAPO Group Size - - - 8
DAPO Epsilon - - - (0.2, 0.3)
Initial KL Coefficient - - - 1e-3
Appendix F The Analysis of ChemDFM-R’s Rationale